If for a two-dimensional potential flow, the velocity potential is given by ϕ = 4x(3y - 4), determine the velocity at the point (2, 3) . Determine also the value of stream function ψ at the poins (2, 3).


1
Expert's answer
2022-02-15T08:50:01-0500

Velocity potential Function

"u = - \frac{{\partial \phi }}{{\partial x}};\;v = - \frac{{\partial \phi }}{{\partial y}}"

Stream Function

"u = \frac{{\partial ψ }}{{\partial y}};\;v = -\frac{{\partial ψ }}{{\partial x}}"


"\u03d5 = 4x(3y - 4)"


"u=-\frac{\delta \phi}{\delta x}" and "v=-\frac{\delta \phi}{\delta y}=-\frac{\delta ψ}{\delta x}"


"u = -4 (3y - 4) =\frac{\delta ψ}{\delta y}"


"\u03c8y = 16y - 6y2 + C1"


and


"v = -12x =-\frac{\delta ψ}{\delta x}"


"\u03c8x = 6x2 + C2"


Therefore, "\u03c8 = \u03c8x + \u03c8y"


"\u03c8 = 6(x2 \u2013 y2) + 16y"


"{\left. ψ \right|_{\left( {2,3} \right)}} = 6\left( {4 - 9} \right) + 16 \times 3"


"= 18" units